OpenStudy (anonymous):

Simplify. Square root -36 plz help!

OpenStudy (anonymous):

6i

OpenStudy (anonymous):

Thank you!

Parth (parthkohli):

\(\color{Black}{\Rightarrow \sqrt{-36} = \sqrt{-1} \times \sqrt{36} = i \times 6 = 6i }\)

OpenStudy (anonymous):

6i in Imaginary Numbers or negative -6

OpenStudy (anonymous):

Thnx everyone!

Parth (parthkohli):

I got that \(\sqrt{-1} \times \sqrt{36}\) because \(\sqrt{ab} = \sqrt a \times \sqrt b\)